OpenStudy (anonymous):

5 x (a+40)= 6 x (a-40) How would you solve this?

OpenStudy (anonymous):

Looks like a distibutive property should be used....

OpenStudy (anonymous):

5a + 200 = 6a - 240 440 = a

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

\[ 5\times(a+40)=6\times(a-40)\\ 5a+200=6a-240\\ -a=-440\\ a=440 \]

OpenStudy (anonymous):

Thank you!